OpenStudy (anonymous):

Hi , every one. I need help to solve these calculus problems, the question is: 1-Find the area under the given curves: a) y = 6x^3 + 6x^2 - 6x between x = 1 and x = 3 b) f(x) = -3e^-3x -3X between x = 0 and x = 2 and other question: Solve the following differential equations: a) dy/dx=3x^2 b) dy/dx=3e^3x c) dy/dx=3xy c) dy/dx=1/(y+1)^2 x e^3x With my thanks to any try..... Adam12

OpenStudy (anonymous):

\[\int\limits_{0}^{2}-3e ^{-3x}-3x = [e ^{-3x}-(3/2)x ^{2} ] _{0}^{2}\]

OpenStudy (anonymous):

then plug in the values and solve

OpenStudy (anonymous):

\[A = \int\limits_{1}^{3} y dx\] \[A = \int\limits_{1}^{3} [6x ^{3} + 6x ^{2} - 6x] dx\] \[\left[ 1.5 x ^{4}+2x ^{3}-3x ^{2}\right]_{1}^{3}\]

OpenStudy (anonymous):

here's the clue for differential \[[dy/dx= ax ^{n} ] = nax^{n-1} , [dy/dx=ae^{nx}] = nae^{nx}\]

OpenStudy (anonymous):

Thank you very much, I rally appreciate your effort answering these problems.

OpenStudy (anonymous):

you're welcome.. if u have any problems again message me, i'm ready to help ..